# EEG-Language Modeling for Pathology Detection

### v3

@@ 4.1 Pretraining comparisons
Table 2: Pathology detection via zero-shot (ZS) and linear probing at 1%, 10%, and 100% labeled data of the TUAB training set. The (second) best scores are printed (underlined) bold. Standard deviations over five model training runs are included.
@@
SV - 71.36$\pm$1.10 81.06$\pm$0.30 84.13$\pm$0.29 - 79.87$\pm$1.30 89.23$\pm$0.51 91.83$\pm$0.32
BYOL - 72.69$\pm$0.57 79.03$\pm$1.16 79.94$\pm$2.14 - 78.85$\pm$0.81 86.75$\pm$0.76 88.82$\pm$0.70
VICReg - 71.76$\pm$0.81 79.6$\pm$1.07 82.46$\pm$0.96 - 78.7$\pm$1.11 86.04$\pm$0.80 88.78$\pm$1.04
ContraWR - 73.30$\pm$1.44 80.72$\pm$1.69 82.44$\pm$1.22 - 80.30$\pm$1.91 86.67$\pm$1.32 88.44$\pm$1.20
RP - 74.52$\pm$1.06 82.16$\pm$0.38 83.36$\pm$0.42 - 82.63$\pm$0.87 89.78$\pm$0.43 91.43$\pm$0.34
TS - 74.99$\pm$0.86 82.16$\pm$0.64 84.10$\pm$0.66 - 82.51$\pm$0.91 89.58$\pm$0.55 91.50$\pm$0.32
CPC - 73.20$\pm$0.79 78.44$\pm$1.00 79.95$\pm$1.49 - 81.48$\pm$1.02 86.44$\pm$1.07 87.92$\pm$1.14
ELM-MIL $l|e$ 68.86$\pm$7.89 82.53$\pm$1.80 86.38$\pm$0.77 87.62$\pm$0.43 75.23$\pm$9.28 89.88$\pm$1.47 92.92$\pm$0.54 93.52$\pm$0.34
ELM-MIL $e|l$ 79.10$\pm$2.93 83.71$\pm$0.59 84.37$\pm$0.97 85.65$\pm$0.97 87.26$\pm$3.19 92.37$\pm$0.43 93.25$\pm$0.27 93.65$\pm$0.16
ELM-MIL $e,l$ 84.31$\pm$0.57 83.10$\pm$0.56 84.21$\pm$0.82 87.11$\pm$0.76 91.56$\pm$1.31 91.54$\pm$0.44 93.14$\pm$0.24 93.91$\pm$0.17
@@
BYOL 63.78$\pm$1.70 76.48$\pm$2.10 80.65$\pm$2.50
ContraWR 65.72$\pm$1.01 72.47$\pm$0.95 75.42$\pm$1.01
VICReg 61.57$\pm$1.49 74.19$\pm$0.63 78.50$\pm$1.60
TS 64.90$\pm$0.70 81.36$\pm$1.53 87.08$\pm$1.02
RP 64.92$\pm$0.81 80.42$\pm$1.83 86.50$\pm$2.17
CPC 65.24$\pm$2.06 77.84$\pm$1.12 79.98$\pm$1.60
ELM-MIL 69.49$\pm$2.26 81.42$\pm$1.15 89.77$\pm$0.21
@@
SPaRCNet Y 0.79M 78.96$\pm$0.18 86.76$\pm$0.12
ContraWR Y 1.6M 77.46$\pm$0.41 84.56$\pm$0.74
CNN-Transformer Y 3.2M 77.77$\pm$0.22 84.61$\pm$0.13
FFCL Y 2.4M 78.48$\pm$0.38 85.69$\pm$0.51
ST-Transformer Y 3.5M 79.66$\pm$0.23 87.07$\pm$0.19
BIOT Y 3.2M 79.59$\pm$0.57 88.15$\pm$0.43
CEReBrO Y 3.58M 79.40$\pm$0.19 87.49$\pm$0.33
CEReBrO Y 40.0M 81.29$\pm$0.15 88.67$\pm$0.06
CEReBrO Y 85.2M 81.67$\pm$0.23 89.16$\pm$0.38
LaBraM-Base Y 5.8M 81.40$\pm$0.19 90.22$\pm$0.09
LaBraM-Large Y 46M 82.26$\pm$0.15 91.27$\pm$0.05
LaBraM-Huge Y 369M 82.58$\pm$0.11 91.62$\pm$0.16
ELM-MIL N 0.93M 84.42$\pm$0.21 91.44$\pm$0.11
Pathology detection. Next, we study the learned representations in their relevance to clinical pathology by training linear probes on TUAB (Figure 2C). Here we observe similar patterns for recording-level classification, with bidirectional ELM-MIL scoring best. Interestingly, good classification was possible with many of the models, causing us to investigate whether the strategy of sub-unit multimodal modeling provides inherent benefits. We provide this additional set of analyses in Appendix A.2, which indicates that our sub-unit alignment strategy promotes the encoding of between-subject information even in the absence of semantically relevant text. This allows ELMe,l to nearly match the best EEG-only pretraining strategy for pathology detection when reports are randomly shuffled.
Table 5: Linear probing for seizure and event classification on the TUSZ and TUEV datasets respectively. The (second) best SSL scores are printed (underlined) bold. Standard deviations over runs are included.
AUROC TUSZ TUEV
Method 1% 10% 100% 100%
SV 84.00$\pm$0.90 87.33$\pm$0.93 90.38$\pm$0.39 86.28$\pm$0.78
BYOL 77.45$\pm$3.14 86.60$\pm$0.79 88.89$\pm$1.02 82.40$\pm$1.60
ContraWR 68.86$\pm$3.13 82.75$\pm$1.66 85.68$\pm$0.74 84.11$\pm$1.64
VICReg 69.61$\pm$1.82 82.01$\pm$0.93 86.17$\pm$0.98 83.26$\pm$1.83
TS 78.60$\pm$1.90 87.63$\pm$0.58 89.77$\pm$0.72 84.86$\pm$1.32
RP 64.41$\pm$2.50 76.72$\pm$1.37 79.52$\pm$1.26 78.95$\pm$1.59
CPC 72.01$\pm$1.00 81.77$\pm$2.24 85.91$\pm$2.00 81.94$\pm$1.75
ELM-MIL 78.98$\pm$5.18 88.98$\pm$0.86 91.51$\pm$0.33 87.69$\pm$1.01
Next, we investigate the unique ability of multimodal language modeling to leverage the language modality to perform ‘zero-shot’ classification (Figure 2D). Without any explicit labels for downstream training, EEG may be classified by computing its similarity in latent space to text prompts representing the candidate classes. As suggested by [12], we create a prompt ensemble over 21 variations of the phrasing ”The EEG is normal, abnormal” (Appendix D). Despite a small dataset, EEG-language models can reach high levels of zero-shot pathology detection if: 1) they include a text projector, 2) include at least the clinician interpretation, and 3) do bidirectional alignment. These models, especially ELM-MILe,l, score highly consistent across different weight initializations, suggesting sound alignment between modalities. Pretraining using exclusively one of the other text clusters yielded poor or inconsistent performance, which follows from these models not being exposed to the explicit phrasing indicating the EEG status as normal or abnormal per se. Their capability likely can be improved by designing appropriate prompts. Having identified the requirements for stable alignment, we focus on the ELM-MILe,l model for EEG-only baselines due to its consistent performance across tasks and seeds.
@@
4.2 Baseline comparisons
4.2.1 Abnormality classification
We compare our ELM-MIL models on the TUAB dataset to six EEG-only methods (Table 2). We pretrain using these methods on the same pretraining dataset as ELMs and, together with the supervised baseline, also an identical EEG encoder architecture. This enabled accurate inference on the effectiveness of the pretraining strategy per se. We find EEG-language pretraining yields large improvements for pathology detection over EEG-only pretraining, with multimodal models being particularly effective at small sample sizes: at 1% of exposed labels, performance increases reach 8.7% balanced accuracy and 9.7% AUROC. To validate our results out-of-distribution, we evaluate learned representations on the NMT dataset without finetuning (Table 3). We find ELM-MILe,l (henceforth ’ELM-MIL’) to still perform well, especially with few labels, where they yield an improvement of at least 3.8% AUROC.
Next, we further contextualize our results in the literature. By probing ELM-MIL representations for 10 second crop-level abnormality classification on TUAB with 80% labels we are able to compare to previously reported evaluations [20, 11, 47]. However, we note that methods use significantly different pretraining data and architectures, which complicates interpretation. Nevertheless, compared to supervised and self-supervised methods, ELM-MIL scores highest and improves over LaBraM-Huge +0.83% on average across scores (Table 4). This is despite LaBraM-Huge requiring finetuning, being pretrained on a large number of datasets (including TUEG), and featuring 369M parameters compared to 0.9M of our EEG encoder. This is likely due to the clinical specificity of ELMs, enabling much smaller models that do not need to be finetuned. As a result, our methodology is also significantly cheaper and faster to train (50 epochs in under 12 hours with 24GB of memory).
@@ 4.2.2 Seizure and event classification
We additionally investigated classification performance based on single, five second EEG crops. Although recording-level classification was the focus of our work, we find ELM to score better than the investigated EEG-only methods on both TUSZ and TUEV as well (Table 5). Due to the smaller size of TUEV, we evaluate only using 100% labels. We provide per-class performance in Appendix A.1.
@@ 4.2.3 Further Ablation Analyses
Whereas for InfoNCE the temperature parameter sets the relative focus across negative samples [48], for MIL-InfoNCE it does so too for positive samples. We therefore test the sensitivity of our methods to the parameter (Figure 3). We find that MIL-InfoNCE is more robust to changes of $\tau$ for pathology detection, while retrieval performance can be further improved by lowering $\tau$. This may be explained as retrieval being subject-based rather than class-based (see Appendix 12). Moreover, performance increases from $\tau<1$ indicates the utility of this additional hyperparameter of InfoNCE, which is absent in NCE.
@@
We perform additional ablations to investigate crucial aspects of the ELM-MIL $e,l$ model. First, we find that additional positive EEG and text samples improve downstream performance (Table 6). We additionally ablate the aggregation method for positive samples and find MIL-InfoNCE to outperform considered alternatives. We compare to aligning only the most similar positive sample (denoted Max+InfoNCE), using attention to create a weighted average across positive samples based on similarity values (Attn+InfoNCE; [49]), as well as taking the sum instead of mean across log-probabilities (Sum+InfoNCE). The latter does not account for the varying amount of text and EEG crops across subjects.
@@ 5 Discussion
The current work presents a first application of multimodal pretraining combining natural language and functional brain data in a medical context. The proposed methodology achieves significantly improved clinical representations compared to EEG-only SSL, while being inexpensive to pretrain. This improvement stems from our novel sub-unit alignment approach in combination with MIL-InfoNCE to address inherent data misalignment challenges. Notably, these multimodal models demonstrated capability for zero-shot pathology detection and label-efficient linear probing with improvements up to 9.7%, an important advancement for clinical applications. We additionally show generalization of ELMs via external validation and clinical event detection tasks.
Some considerations of this study deserve mention. The current limitation of publicly available paired EEG-report datasets presents a challenge for scaling pretraining data. Future work could address this through the generation of synthetic text captions based on clinical metadata. Furthermore, due to computational constraints, analyses on scaling model sizes are yet to be performed. However, we find highly favourable performance of ELMs compared to significantly larger EEG foundation models pretrained on a large number of datasets.
6 Acknowledgments and Disclosure of Funding
This research was funded by the Deutsche Forschungsgemeinschaft (DFG) through FOR 5187 (project number 442075332). Additional support was provided by the DFG through the following projects: CRC 1404 (project number 414984028), TRR 265 (project number 402170461), and RU 5363 (project number 459422098).
The data used in this study was provided by the Neural Engineering Data Consortium at Temple University. For further details about this data, please access the following URL: https://isip.piconepress.com/projects/tuh_eeg/html/downloads.shtml.
@@ Appendix A Additional Results
A.1 TUEV: Per-Class analysis
@@
A.2 Language-independent effects of sub-unit alignment.
@@
A.3 Post-hoc investigation of data leakage
@@ B.1 Optimization
All models are pretrained using the LARS optimizer [50] with a cosine decay learning rate schedule over 50 epochs, with a warm-up of 4 epochs. The base learning rate is set to 0.3 for EEG-only, 0.01 for ELMs, and 0.06 for ELM-MIL, scaled with the batch size (BaseLR $\times$ BatchSize/256; [41]). We use a weight-decay parameter of 1 $\times$ 10-4. Models were trained on either an Nvidia Geforce GTX 3090 or Tesla V100 GPU and require less than 24GB of memory. Training took approximately 9 hours for EEG-language modeling or 18 hours for EEG-only modeling due to data augmentations. We used CUDA v11.3 and PyTorch v1.12.1.
@@ B.2 EEG Encoder
We use a CNN architecture with a residual stream as the EEG encoder for all analyses (Figure 6). The model uses parallel convolutions, involving reflection padding and 1D-convolutions with kernel sizes $\{4,8,16\}$ with 32 filters each. These outputs are concatenated, resulting in a 96 dimensional representation and 747K trainable parameters. We compare input lengths of EEG crops varying from 5 to 60 seconds. This presents a trade-off where longer crops result in a greater information content per crop, while reducing the total sample size. As EEG-only pretraining relies on data augmentations, this introduces an additional influence of crop length. Specifically, longer crop lengths likely make the pretraining task easier, as augmentations introduce relatively lesser distortion due to the greater information content. We therefore compare performance of different crop lengths for both EEG-language and EEG-only pretraining. As the EEG encoder progressively downsamples the signal, we adjust the pooling layers to the input length. These adjustments are shown in Table 8. For EEG-language pretraining we evaluate zero-shot pathology detection, while for EEG-only pretraining we are required to compare the performance of a linear probe. Results are shown in Figure 7. Due to computational resources, we only compare crop lengths for BYOL and ELMl as representations of EEG-only and EEG-language modeling. We observe that for EEG-only pretraining an intermediate crop-length of 20 seconds performs best, which matches the findings by [8]. Meanwhile, zero-shot pathology detection is found to be relatively insensitive to crop lengths of at least 10 seconds, with 60 second crops scoring highest, while the shortest crop length showed unstable learning. For TUSZ and TUEV evaluations, this was solved by lowering the learning rate.

### v4

@@ 4.1 Pretraining comparisons
Table 2: Pathology detection via zero-shot (ZS) and linear probing at 1%, 10%, and 100% labeled data of the TUAB training set. The (second) best scores are printed (underlined) bold. Standard deviations over five model training runs are included. Supervised serves as a reference and refers to training end-to-end directly on labels.
@@
[…]
Clinical phenotyping. We study the learned representations in their relevance to clinical pathology by training linear probes on TUAB (Figure 2C). Here we observe similar patterns for recording-level classification, with bidirectional ELM-MIL scoring best. Interestingly, good classification was possible with many of the models, causing us to investigate whether the strategy of sub-unit multimodal modeling provides inherent benefits. We provide this additional set of analyses in Appendix A.3, which indicates that our sub-unit alignment strategy promotes the encoding of between-subject information even in the absence of semantically relevant text. This allows ELMe,l to nearly match the best EEG-only pretraining strategy for pathology detection when reports are randomly shuffled.
@@
[…]
Next, we investigate the unique ability of multimodal language modeling to leverage the language modality to perform ‘zero-shot’ classification (Figure 2D). Without any explicit labels for downstream training, EEG may be classified by computing its similarity in latent space to text prompts representing the candidate classes. As suggested by Radford et al. (2021), we create a prompt ensemble over 21 variations of the phrasing ”The EEG is normal, abnormal” (Appendix D). Despite a small dataset, EEG-language models can reach high levels of zero-shot pathology detection if: 1) they include a text projector, 2) include at least the clinician interpretation, and 3) do bidirectional alignment. These models, especially ELM-MILe,l, score highly consistent across different weight initializations, suggesting sound alignment between modalities. Pretraining using exclusively one of the other text clusters yielded poor performance, which follows from these models not being exposed to the explicit phrasing indicating the EEG status as normal or abnormal per se. Their capability likely can be improved by designing appropriate prompts. Having identified the requirements for stable alignment, we focus on the ELM-MILe,l model for EEG-only baseline comparisons due to its consistent performance across tasks and seeds.
4.2 Baseline comparisons
4.2.1 Abnormality classification
We compare our ELM-MIL models on the TUAB dataset to six EEG-only methods (Table 2). We pretrain using these methods on the same pretraining dataset as ELMs and, together with the supervised baseline, also an identical EEG encoder architecture. This enabled accurate inference on the effectiveness of the pretraining strategy per se. We find ELMs yield large improvements for pathology detection over EEG-only pretraining, with multimodal models being particularly effective at small sample sizes: at 1% of exposed labels, performance increases reach 8.7% balanced accuracy and 9.7% AUROC. To validate our results out-of-distribution, we evaluate learned representations on the NMT dataset without finetuning (Table 3). We find ELM-MILe,l (henceforth ’ELM-MIL’) to still perform well, especially with few labels, where they yield an improvement of at least 3.8% AUROC.
@@ 4.2.1 Abnormality classification
[…]
@@
Table 6: Linear probing for seizure and event classification on the TUSZ and TUEV datasets respectively.
AUROC TUSZ TUEV
Method 1% 10% 100% 100%
[…]
Next, we further contextualize our results in the literature. By probing ELM-MIL representations for 10 second crop-level abnormality classification on TUAB with 80% labels we are able to compare to previously reported evaluations (Yang et al., 2024; Jiang et al., 2024; Dimofte et al., 2025). However, we note that methods use significantly different pretraining data and architectures, which complicates interpretation. Nevertheless, ELM-MIL scores highest and improves over LaBraM-Huge +0.83% on average across scores (Table 4). This is despite LaBraM-Huge requiring finetuning, being pretrained on a large number of datasets (including TUEG), and featuring 369M parameters compared to 0.9M of our EEG encoder. To further compare the quality of learned representations, we apply our linear probing pipeline to the LaBraM-Base model (Appendix 7-9). We find gains for our ELM of up to 5.7–13.4% depending on the dataset, highlighting the improved semantic content in ELM representations. This stark discrepancy is likely due to the clinical specificity of ELMs, enabling much smaller models that do not need to be finetuned. As a result, our methodology is also significantly cheaper and faster to train (50 epochs in under 12 hours with 24GB of memory).
@@ 4.2.2 Seizure and event classification
We additionally investigated classification performance based on single, five second EEG crops. Although recording-level classification was the focus of our work, we find ELM to score better than the investigated EEG-only methods also on both seizure detection (TUSZ) and 6-class event detection (TUEV; Table 6). We provide per-class performance in Appendix A.2.
@@ 4.2.3 Further Ablation Analyses
Whereas for InfoNCE the temperature parameter sets the relative focus across negative samples (Wang & Liu, 2021), for MIL-InfoNCE it does so too for positive samples. We therefore test the sensitivity of our methods to the parameter (Figure 3). We find that MIL-InfoNCE is more robust to changes of $\tau$ for pathology detection, while retrieval performance can be further improved by lowering $\tau$. This may be explained as retrieval being subject-based rather than class-based (see Appendix 13). Moreover, performance increases from $\tau<1$ indicates the utility of this additional hyperparameter of InfoNCE, which is absent in NCE.
@@
We perform additional ablations to investigate crucial aspects of the ELM-MIL $e,l$ model. First, we find that additional positive EEG and text samples improve downstream performance (Table 5). We additionally ablate the aggregation method for positive samples and find MIL-InfoNCE to outperform considered alternatives. We compare to aligning only the most similar positive sample (denoted Max+InfoNCE), using attention to create a weighted average across positive samples based on similarity values (Attn+InfoNCE; (Ilse et al., 2018)), as well as taking the sum instead of mean across log-probabilities (Sum+InfoNCE). The latter does not account for the varying amount of text and EEG crops across subjects.
4.2.4 Alignment visualizations
For model interpretability, we visualize temporal multimodal alignment in example hold-out recordings (Figure 4). We compute cosine similarity between all EEG crop embeddings in a recording and an embedded snippet from its paired clinical report, then plot crops with the highest and lowest similarity. We observe distinct periods of stronger/weaker alignment, where strong alignment corresponds to relevant clinical events. This indicates our ELMs achieve temporally selective alignment of clinical EEG events without explicit temporal event information or labels, consistently across independent pretraining runs. Additional examples, including failure cases, are in Appendix E.3.
Figure 4: Left, top: The cosine similarity between EEG embeddings of 5-second crops and the embedding of the paired text ”seizures arising from the right hemisphere” shows distinct periods of stronger and weaker alignment. Shading indicates the standard deviation across 5 random weight initializations. Green and purple dots mark the two crops with the highest and lowest similarity respectively. Left, bottom: EEG traces of the crops indicated by the coloured dots. Electrodes are colored based on their position (left: red, right: blue) and clearly reveal right-lateral seizure activity. Right: As left-sided plot, but for a different recording with the text ”poly spike and slow wake”.
@@ 5 Discussion
This paper presents a first application of multimodal pretraining combining natural language and functional brain data in a medical context. The proposed methodology achieves significantly improved clinical representations compared to EEG-only SSL, while being inexpensive to pretrain. This improvement stems from our novel sub-unit alignment approach in combination with MIL-InfoNCE to address inherent data misalignment challenges. Notably, these multimodal models enabled zero-shot pathology detection and label-efficient linear probing with improvements up to 9.7%. We additionally show generalization of ELMs via external validation and clinical event detection tasks. Our pathology-sensitive multimodal alignment is a critical step toward automated report generation (e.g. Biswal et al. 2020), ensuring EEG-text representations capture clinical information for future documentation tasks.
Some considerations of this study deserve mention. The current limitation of publicly available paired EEG-report datasets presents a challenge for scaling pretraining data. Future work could address this through the generation of synthetic text captions based on clinical metadata. Furthermore, due to computational constraints, analyses on scaling model sizes are yet to be performed. However, we find highly favourable performance of ELMs compared to significantly larger EEG foundation models pretrained on a large number of datasets. We provide code and pretrained models at https://github.com/SamGijsen/ELM.
Acknowledgements
This research was funded by the Deutsche Forschungs-gemeinschaft (DFG) through FOR 5187 (project number 442075332). Additional support was provided by the DFG through the following projects: CRC 1404 (project number 414984028), TRR 265 (project number 402170461), and RU 5363 (project number 459422098).
The data used in this study was provided by the Neural Engineering Data Consortium at Temple University. For further details about this data, please access the following URL: https://isip.piconepress.com/projects/tuh_eeg/html/.
@@ Acknowledgements
Impact Statement
Performance across both recording-level classification and event detection tasks suggests that our model learns clinically relevant features at multiple temporal scales. With further progress, this capability may support various future clinical applications, from rapid screening of prolonged recordings to real-time event detection. Whereas the present study focuses on establishing an initial application to explore viability, future work may benefit from focusing on improving the interpretability of these representations through techniques such as channel-specific attribution. The multimodal nature of our approach, by aligning EEG with clinical reports in a pathology-sensitive manner, not only enhances detection but also lays an important foundation for automated report generation. Specifically, such generation may greatly benefit from an aligned latent space which contains clinical information. This could facilitate clinical documentation by translating EEG signals into structured summaries. These can constitute highly valuable future efforts given the time-intensive nature of manual reporting. Certain clinical limitations also deserve further attention, such as a careful study of how the frequency of specific pathology and clinical events in reports impacts model performance. Finally, biases present in language models may impact multimodal pretraining, which should be investigated in future work.
@@ Appendix A Additional Results
A.1 Linear probing comparisons with LaBraM
We apply our linear probing pipeline to the LaBraM-Base model and find significant performance drops, highlighting its requirement for downstream finetuning and the improved clinical representations of ELMs. We perform these analyses for NMT (Table 7), TUAB (Table 8), and TUEV (Table 9), while omitting TUSZ as this data was included in LaBraM pretraining. Standard deviations indicate variability across five repetitions of 10-fold cross-validation. Whereas for ELM-MIL we use a model trained with different weight initialization for each repetition, for LaBraM only one pretrained model is made available which we use for every repetition.
Table 7: Linear probing for abnormality classification comparing ELM-MIL and LaBraM on the NMT dataset using 1%, 10%, and 100% labeled training data.
Balanced Accuracy AUROC
Method 1% 10% 100% 1% 10% 100%
[…]
Table 8: Linear probing for abnormality classification comparing ELM-MIL and LaBraM on the TUAB dataset using Zero-Shot, 1%, 10%, and 100% labeled training data.
Balanced Accuracy AUROC
Method ZS 1% 10% 100% ZS 1% 10% 100%
[…]
Table 9: Linear probing for abnormality classification comparing ELM-MIL and LaBraM on the TUEV dataset using 80% labeled training data.
BACC AUROC
Method 80% 80%
[…]
A.2 TUEV: Per-Class analysis
@@
A.3 Language-independent effects of sub-unit alignment.
@@
A.4 Post-hoc investigation of data leakage
@@ B.1 Optimization
All models are pretrained using the LARS optimizer (You et al., 2017) with a cosine decay learning rate schedule over 50 epochs, with a warm-up of 4 epochs. The base learning rate is set to 0.3 for EEG-only, 0.01 for ELMs, and 0.06 for ELM-MIL, scaled with the batch size (BaseLR $\times$ BatchSize/256; (Grill et al., 2020)). When training on 5-second crops, we lower the learning rate for EEG-only to 0.1 and ELM-MIL to 0.02 to avoid instability. We use a weight-decay parameter of 1 $\times$ 10-4. Models were trained on either an Nvidia Geforce GTX 3090 or Tesla V100 GPU and require less than 24GB of memory. Training took approximately 9 hours for EEG-language modeling or 18 hours for EEG-only modeling due to data augmentations. We used CUDA v11.3 and PyTorch v1.12.1.
@@ B.2 EEG Encoder
We use a CNN architecture with a residual stream as the EEG encoder for all analyses (Figure 7). The model uses parallel convolutions, involving reflection padding and 1D-convolutions with kernel sizes $\{4,8,16\}$ with 32 filters each. These outputs are concatenated, resulting in a 96 dimensional representation and 747K trainable parameters. We compare input lengths of EEG crops varying from 5 to 60 seconds. This presents a trade-off where longer crops result in a greater information content per crop, while reducing the total sample size. As EEG-only pretraining relies on data augmentations, this introduces an additional influence of crop length. Specifically, longer crop lengths likely make the pretraining task easier, as augmentations introduce relatively lesser distortion due to the greater information content. We therefore compare performance of different crop lengths for both EEG-language and EEG-only pretraining. As the EEG encoder progressively downsamples the signal, we adjust the pooling layers to the input length. These adjustments are shown in Table 11. For EEG-language pretraining we evaluate zero-shot pathology detection, while for EEG-only pretraining we are required to compare the performance of a linear probe. Results are shown in Figure 8. Due to computational resources, we only compare crop lengths for BYOL and ELMl as representations of EEG-only and EEG-language modeling. We observe that for EEG-only pretraining an intermediate crop-length of 20 seconds performs best, which matches the findings by Mohsenvand et al. (2020). Meanwhile, zero-shot pathology detection is found to be relatively insensitive to crop lengths of at least 10 seconds, with 60 second crops scoring highest, while the shortest crop length showed unstable learning. For TUSZ and TUEV evaluations, this was solved by lowering the learning rate.